OpenStudy (vera_ewing):

What logarithmic function represents the data in the table? x f(x) 216 3 1296 4 7776 5 (2 points) f(x) = logx 5 f(x) = 6 log x f(x) = log 6x f(x) = log6 x

OpenStudy (vera_ewing):

@zepdrix

OpenStudy (dugalde):

go back to the definition. what is a log? well, it is the value you get when you express a number to a certain power. The easy way to recomember is: log to the base 10 of 100 is 2, because 10^2 = 100 now, what number exists such that: x^3 = 216 x^4 = 1296 x^5 = 7776 and will fit that pattern? well, 6^3 = 216 and, 10^2 = 100 just replay the words. The nice thing about mathematics is that athe process does not change just because the numbers do.